# Evidence for Low Universal Equilibrium Black Hole Spin in Luminous Magnetically Arrested Disks

### v1

@@ 4.3 Observational implications
BL acknowledges support by a National Science Foundation Graduate Research Fellowship under Grant No. DGE-2234667. BL also acknowledges support by a Illinois Space Grant Consortium (ISGC) Graduate Fellowship supported by a National Aeronautics and Space Administration (NASA) grant awarded to the ISGC. JJ acknowledges support by the NSF AST-2009884, NASA 80NSSC21K1746 and NASA XMM-Newton 80NSSC22K0799 grants. AT acknowledges support by NASA 80NSSC22K0031, 80NSSC22K0799, 80NSSC18K0565 and 80NSSC21K1746 grants, and by the NSF grants AST-2009884, AST-2107839, AST-1815304, AST-1911080, AST-2206471, AST-2407475, OAC-2031997. This work was performed in part at the Kavli Institute for Theoretical Physics (KITP) supported by grant NSF PHY-2309135. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC allocations m4603 (award NP-ERCAP0029085) and m2401. The computations in this work were, in part, run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation. An award of computer time was provided by the ASCR Leadership Computing Challenge (ALCC), Innovative and Novel Computational Impact on Theory and Experiment (INCITE), and OLCF Director’s Discretionary Allocation programs under award PHY129. This research was partially carried out using resources from Calcul Quebec (http://www.calculquebec.ca) and Compute Canada (http://www.computecanada.ca) under RAPI xsp-772-ab (PI: Daryl Haggard). This research also used HPC and visualization resources provided by the Texas Advanced Computing Center (TACC) at The University of Texas at Austin, which contributed to our results via the LRAC allocation AST20011 (http://www.tacc. utexas.edu).
@@ Appendix A Convergence test
We compare low- and high-resolution simulations for disk scale heights of $h/r=0.05$ and $h/r=0.1$. Figure 13 shows spin-up parameter values, with the primary result simulation, at standard low resolution for $h/r=0.1$, shown in purple, yielding an equilibrium spin of $0.29$. To verify that the disk scale height is fully resolved, we ran a higher-resolution simulation on OLCF’s Summit (red in Figure 13). The consistency of $s$ between resolutions confirms that lower-resolution simulations are sufficient for $h/r=0.1$ and are therefore used for analysis in the main text.
We include lower-resolution simulations with $h/r=0.05$ (blue) for reference, though they are not used for analysis, while the high-resolution simulation is shown in green. For a lower thermal scale height, $h/r=0.05$, we observe significant differences in $s$ between low-resolution simulations run on a workstation and high-resolution simulations on Summit. Figure 13 shows that low-resolution simulations systematically overestimate $s$, leading to an equilibrium spin overestimate of about 30%. Analyzing individual contributions to $s$ (not shown) reveals that this discrepancy arises from an overestimation of hydrodynamic energy and angular momentum fluxes. We suspect that low-resolution $h/r=0.05$ simulations fail to fully resolve the fastest-growing MRI mode, reducing turbulent dissipation, which in turn impacts these fluxes. For this reason we choose to not use the low resolution $h/r=0.05$ for analysis in the main text.
We conclude that higher resolutions are necessary for MADs with $h/r\leq 0.05$ to accurately determine the equilibrium spin.
Figure 13: Convergence test for two disk scale heights of $h/r=0.1$ and $0.05$. Purple and blue circles show low resolution (LR) simulations (give those from table). Red and green crosses show high resolution (HR) simulations run on Summit, with the same parameters as the LR simulations. The plot only shows results for spins between $-0.1$ and $0.6$ to highlight the convergence near the equilibrium spins.

### v2

@@ Acknowledgements.
BL acknowledges support by a National Science Foundation (NSF) Graduate Research Fellowship under Grant No. DGE-2234667. BL also acknowledges support by a Illinois Space Grant Consortium (ISGC) Graduate Fellowship supported by a National Aeronautics and Space Administration (NASA) grant awarded to the ISGC. JJ acknowledges support by the NSF AST-2009884, NASA 80NSSC21K1746 and NASA XMM-Newton 80NSSC22K0799 grants. JJ also acknowledges support from NASA 80NSSC22K0826 and NSF AST 2307983. AT acknowledges support by NASA 80NSSC22K0031, 80NSSC22K0799, 80NSSC18K0565 and 80NSSC21K1746 grants, and by the NSF grants AST-2009884, AST-2107839, AST-1815304, AST-1911080, AST-2206471, AST-2407475, AST-2510570, OAC-2031997. This work was performed in part at the Kavli Institute for Theoretical Physics (KITP) supported by grant NSF PHY-2309135. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC allocations m4603 (award NP-ERCAP0029085) and m2401. The computations in this work were, in part, run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation. An award of computer time was provided by the ASCR Leadership Computing Challenge (ALCC), Innovative and Novel Computational Impact on Theory and Experiment (INCITE), and OLCF Director’s Discretionary Allocation programs under award PHY129. This research was partially carried out using resources from Calcul Quebec (http://www.calculquebec.ca) and Compute Canada (http://www.computecanada.ca) under RAPI xsp-772-ab (PI: Daryl Haggard). This research also used HPC and visualization resources provided by the Texas Advanced Computing Center (TACC) at The University of Texas at Austin, which contributed to our results via the LRAC allocation AST20011 (http://www.tacc. utexas.edu).
@@ Appendix A Convergence test
We compare low- and high-resolution simulations for MAD scale heights of $h/r=0.05$ and $h/r=0.1$. Figure 13 shows in purple the spin-up parameter values for our fiducial $h/r=0.1$ simulation suite: it uses the standard (“low”) resolution and yields an equilibrium spin of $a_{\rm eq,MAD}^{h/r=0.1}=0.29$. To verify that these results are converged with the numerical resolution, we ran a higher-resolution simulation on OLCF Summit supercomputer and show its result in red. The consistency of $s$ between the resolutions demonstrates that the lower resolution simulations are sufficient to accurately determine the $s$ values for $h/r=0.1$ MADs and are therefore used for the analysis in the main text.
For thinner, $h/r=0.05$ MADs, Figure 13 reveals significant differences in $s$ between low-resolution simulations (shown in green and sufficiently small to be run on a workstation) and high-resolution simulations (shown in blue and run on OLCF Summit). In fact, low-resolution simulations systematically overestimate the spin-up parameter, $s$ (by $\Delta s\sim 1$), and the equilibrium spin, $a$ (by $\Delta a\sim 0.1$). Analyzing individual contributions to $s$ (not shown) reveals that this discrepancy arises from an overestimation of hydrodynamic energy, $e_{\rm HD}$, and angular momentum, $l_{\rm HD}$, fluxes. This may be due to our low-resolution $h/r=0.05$ simulations not fully resolving the fastest-growing MRI mode, reducing turbulent dissipation and impacting these fluxes.
Summing up, the resolution of $\tilde{N}_{\theta}=5$ (55 cells per disk scale height) is insufficient to resolve thin MADs (with $0.05\lesssim h/r\lesssim 0.1$). To recover the accurate values of $s$, we find that we need $\tilde{N}_{\theta}\simeq 10$. For this reason, we do not use the low-resolution $h/r=0.05$ simulations for the analysis in the main text.
Figure 13: Thin MAD simulations need a vertical resolution of $\tilde{N}_{\theta}\gtrsim 9$ cells per scale height to yield accurate values of the spin-up parameter, as seen in a convergence study for thin MADs (with $h/r=0.1$ and $0.05$). Purple and blue circles show the low-resolution (LR) simulations and red and green crosses show their high-resolution (HR) counterparts (see Table 1). The figure only shows results for spins between 0.1-0.1 and $0.6$ to highlight the convergence near the equilibrium spins.